Balloon-born hard X-ray imaging experiment SUMIT (formerly known as NUSMIT ) performed flight operation in Cachoeira Paulista, SP, Brazil.
Launch is scheduled on Nov. 8, 2006.
Update on Nov. 8, 2006 : Model prediction of float wind indicates that the best condition is available from Nov. 12. We postpone launch, and reschedule on day-to-day basis.
Update on Nov. 11, 2006 : Revised launch dats is Nov. 13.
Update on Nov. 13, 2006 : Launch is postponed until Nov. 17.
Update on Nov. 16, 2006 : Due to bad condition of float wind, launch is postponed until Nov. 19.
Update on Nov. 17, 2006 : We will make a launch attempt tomorrow, Nov. 18.
Update on Nov. 18, 2006 : Launch canceled. We will make a launch attempt tomorrow, Nov. 19.
SUMIT (SUperMirror Imaging Telescope experiment) is a joint program between Nagoya University and Osaka University, also in collaboration between Chubu University, ISAS/JAXA and JASRI/SPring-8 on instrument development. Launch operation is carried out by ISAS/JAXA Balloon Center and INPE Balloon Sector. Also daily technical supports are provided by INPE Balloon Sector during the campaign.
It has been two days after we finished balloon flight. I apologize I could not update this page, because we have been terribly busy for packing.
We verified all system functioning fine during ascent, but we had power outage of ground system because of thunderstorm a little before reaching to float. Just before this we had lost science telemetry. It took us while to restart ground systems and to establish communication with gondola, but we succeeded to restart operation. Unfortunately science telemetry did not recover until very last period of observation. After operation had resumed, we started pointing. Pointing telemetry was fine, and we confirmed that the telescope was tracking observation target, and we also knew that instruments had been working fine during ascent. Therefore we are very confident that observation was successful, although we could not recover science data. Although it was not "loud and clear", science telemetry came back just before termination, so we will try to salvage data.
The float wind headed south unexpectedly. Balloon headed to the ocean. We tried to change course by changing altitude, but we could not get it back. We terminated flight above the Atlantic Ocean, and payload has not been recovered. Last known coordinate of the payload was 43:16:02W and 24:04:00S.
We will finish packing for return shipment today, and leave balloon base this afternoon. We will stay at Sao Paulo tonight, and leave for Japan tomorrow. It will take us about 36 hours to home.
By this update I finish live reports from Brazil. I appreciate your kind cooperation.
Department of Physics, Nagoya University, Japan
We have terminated the flight, and will start descenting in a moment.
Altitude 32km. We confirmed that attitude actuators are functioning normal.
We are at 28km. Focal planes and Star Camera are functioning normal.
The launch was picture perfect. God speed!
We cleared altitude of 10km.
Inflation has started.
Gondola is at launch position. Waiting for inflation.
Balloon is out of the box.
Experiment is powered up. Waiting for launch.
We will turn on the experiment in 30 min.
Arrived at the launch pad.
Picked up by launcher crane. Moving to the launch pad.
Pre-roll-out telemetry check has completed. Starting roll-out procedure.
We start launch procedure. Pick-up time is 14:20.
We've switched to internal battery, and waiting for ISAS's decision.
We make a launch attempt today, although weather forecast for this afternoon is not very promising. GO/NOGO decision will be made at 13:00. Yesterday's launch was scrubbed because weather prediction at launch time was rain. We waited for an hour to get an update of forecast but the prediction did not change.
Due to bad weather prediction launch operation has been canceled. We will make an attempt tomorrow again.
Weather forecast predicts a shower soon. We hold until 15:30.
We finished final telemetry check with internal battery. Holding for roll-out.
We schedule launch today. Yesterday we finished flight preparations for today. This morning ISAS has launched sounding balloon. Ascenting smoothly so far.
Prediction of surface weather of Nov. 19 is bad, so we will make launch attempt tomorrow, Nov. 18. Launch time is 18:00. ISAS will launch sounding balloon in the morning, and we will make a decision at noon.
IMPORTANT NOTICE: Due to power failure and server shutdown, status report during the flight will be unavailable.
Temporal page is prepared.
Launch is postponed, and today is two days before launch. Yesterday we did recovery briefing with ISAS. We also did pointing test and star tracking test. We are waiting for weather, so we do these tests as daily routine.
Latest wind forecast indicates we are still in the turn around, and thus we are unable to launch from this location. Launch is postponed until Nov. 19. This year's condition is unusual, and very difficult to make a firm prediction. We keep rescheduling on day-to-day basis.
One day before launch. However again, wind prediction for tomorrow looks less promising. We schedule weather briefing at 11:30. We took half-day off yesterday, except for those who had some left-over tasks. We did star tracking test in the evening.
Two days before launch. Yesterday we did tracking practice, briefing of recovery procedure, and other misc flight preps.
Launch was scrubbed yesterday, and we retrieved equipments from Telemetry Building and set them up in the hanger.
Latest prediction of float wind indicates bad trajectory. Launch scrubbed. Next launch opportunity is Nov. 17.
We performed Dressed Rehearsal yesterday morning. We schedule launch today. Weather forecast, either of surface and float, is not very promising. We have all-hand-meeting on 14:00 and make a decision. We will lift off at 18:00.
We performed Dressed Rehearsal, and declared flight ready. We will launch at 18:00 tomorrow (Nov. 13), local time. Pick-up is scheduled on 14:20.
We scheduled Final Compatibility Check (or "Dressed Rehearsal", "Launch Pad Compatibility Check", whatever.) We follow all launch procedures except for balloon layout. We started from GSE setup at Telemetry Building, followed by startup-shutdown checkout in hanger with internal battery. But scrubbed just before pick up because of rain. We will do it again tomorrow. From now on all operations are done through radio communication to LASCOS. We did pointing practice in the evening.
No change on launch schedule; Nov. 13.
We continued pointing practice in the morning. We took half-day-off in the afternoon. Launch is scheduled on Nov. 13.
We spent whole day for final preparation and dressing of the experiment. We checked flight-line procedure, and performed rehearsal. We found slight glitch in decoupling motor motion, so replaced it with new one.
We had pizza for dinner but it's not just a pizza. We went down to pizza restaurant and bought pizza "to go". Of course we had no idea how to tell "to go" in Portuguese. We made sentences in Portuguese using Google translation, wrote them down, and showed it to them. Somehow they understood what we want. Unexpected problem was that we had to choose whether we bring pizza back by ourselves, or let them deliver. Obviously we didn't understand what they are asking. Then, the guy at the restaurant brought a pizza box to us, tapped it, and then pointed himself, then us, than swung his arm towards the door. Meaning, "Do I bring pizza, or do you?". Ah! Communication established! We answered "nos", one of very few words we knew, and we got four family-size pizza.
ISAS put ballast today, and total weight was measured as 1525kg (with ballast of 260 kg). The number was about 3% heavier than expected. We also performed telemetry compatibility check with LASCOS again.
We had dinner at "Kanpai Shokudo" (Diner Toast). Ginza was good, but Kanpai was good too. The "best choice of the day" was Fukaya's "Tekka maki", a tuna roll. This place again, miso soup was excellent.
We performed compatibility test today. Command and telemetry compatibility, mechanical fit check, and fine tuning of telemetry receiver have been done. ISAS calculated prediction of trajectory, and found Nov. 12 is the best condition.
According to a model prediction of float wind and a trajectory thus calculated, the condition becomes better from Nov. 12. We keep updating forecast and determine launch date on day-to-day basis. Basically we decide launch date two days in advance to show time. Before Nov. 12 balloon may approach to concentrated area, which definitely must be avoided.
Flight preparation continues. Compatibility test is scheduled tomorrow.
ISAS launched another sounding balloon, BT1, today. Balloon reached 32km altitude as expected. "Partial" compatibility test was done in the afternoon. We are not perfectly compatible yet. Pointing test was don until midnight.
Wiring to ISAS and INPE module continues, as well as flight preparation of our cabling. Late afternoon we calibrated magnetometer outside. INPE/SJ helped us determining geological position of reference landmark.
Dinner was at "Ginza" in Lorena. Miso soup was excellent.
ISAS has launched sounding balloon today. Wiring and connection to ISAS modules continue. We repeat pointing test every day, and parameters are about to be finalized.
Pointing test continues. It has been raining for couple of days, and because of leak on the roof we had to suspend testing when rain is heavy. ISAS is preparing our balloon.
Max elevation check was done today. Interference between mirror bench and rigging structure may occur near maximum elevation. So far we could go up to 75 deg safely.
Not much to say. Pointing test, checking of flight code, and other flight preparations continue. On Oct. 30 parties from ISAS and INPE arrived.
Pointing test continues. Today, upper-frame strut was modified, and interference of wire and frame was solved. We tested star cameras until midnight, and automatic solution was confirmed ok.
Reaction wheel has been fixed, and pointing test was resumed. Star camera test continues too. Another ISAS cargo arrived. So did balloon.
We went downtown Cachoeira to buy some electric equipments. Although we had difficulty in communication, it was fun. But no success. I mean, we could somehow tell them what we need, but they didn't have unfortunately. Anyway we need more Portuguese practice.
We had a mechanical trouble with reaction wheel, and spent whole day for replacing it. Gondola was not picked up today. Weather was not perfect, and Star Camera test was canceled.
Elisete and Oscar arrived from INPE Sao Jose, to set up their equipments.
Strut structure is now mounted, and gondola structure is very close to flight configuration. For next several days, we test pointing function during the day, and test star camera in the night.
LASCOS has arrived from ISAS.
We've moved to another hanger for hanging test. Rest of the days will be spent for pointing tests. Weather was pretty fine, so we tested Star Camera till midnight.
Prof. Tsunemi has left for Japan. He will be back in 10 days.
We had a photo session today.
SD-CCD device was installed, and scientific instruments are (almost) ready to fly, unless something is forgotten. Also focus of Star Camera is adjusted. Pointing test is scheduled to start tomorrow.
We took a day-off. Some of us drove to Paraty.
Alignment of mirrors has been completed. With long focal length of 8m, structural distortion affects alignment roughly twice as large as Suzaku satellite. Currently focus position is aligned within 1 arcmin error, and optical axes with 30 arcsec. They are acceptable numbers, and are also comparable to previous satellite missions such as ASCA and Suzaku. This was the first major milestone, and next is pointing test.
While alignment continues, other flight preparations such as SD-CCD BAT installation, strut frame construction, thermal insulators and crush pads are in progress. In a few days we start pointing test with entire gondola picked up.
Still doing alignment. This time again, reference cameras and mirrors give results slightly different from pre-ship-out tuning. Moreover multiple mirrors are mounted for the first time, and add more complexity. Frustrating. May take a while.
Prof. Kunieda (Nagoya, PI) and Prof. Tsunemi (Osaka) have arrived.
Alignment continues and don't have much to say. It is a boring process. Construction of upper strut and thermal insulators continues.
The most irritating part of the experiment has started; alignment. We are now aligning mirrors and focal plane detectors. SD-CCD cleared communication check. Construction of upper-frame strut has started.
The flag is to visualize the rotation of elevation drive shaft.
Final instrument aboard is SD-CCD, and has been mounted today. CCD chip is not delivered yet, but rest of the system is functioning normal. Mirror front guard was installed. Not to say much about HXI, because it is working very fine. Main instruments are installed, and alignment starts tomorrow.
Installation of science payload continues. Mirrors were mounted today. It was a hot day, and mirror temperature reached close to 35 degC, which is an upper limit of guaranteed temperature. So we constructed cooling system with air conditioner taken off from the wall, and plastic films. It worked.
Installation of HXI (one of focal plane imager) has been completed. Cabling is done, and it is functioning normal. Preparation of mirror is in progress. We will install mirrors tomorrow. Osaka crew has arrived (Emi Miyata and Kazu Ikegami). They are unpacking SD-CCD detector.
We've started installation of science payload. The lid on sensor side of HXI pressure vessel was closed, and will not open again unless there is any trouble. Installation of mirror heater is in progress. We spend next two to three days for final installation. Cargo from Osaka University has arrived. Looks ok.
We took a half-day-off this morning. Installation of small parts and other misc. procedures were done. Today is a national holiday "Nossa Senhora de Aparecida." We received warm messages from InFOCuS people at NASA Goddard. Most of us from Nagoya also works for InFOCuS collaboration, and we are planning to "return to flight" very soon. Many thanks to Takashi, Steve and Doug. Steve, we are following exactly what you taught us how to make crush pads.
Installation onto optical bench has begun. So far star camera and alignment monitor have been installed, and cabling is underway. First star-pointing test was carried out. Preparation of mirrors will start on Friday, and also SD-CCD instruments and crew will arrive on Friday too. We expect their installation is completed on Sunday.
Power-up/down procedure test has been done for instruments on gondola base. All instruments are functioning normal so far. Alignment of base instruments are also done, and analysis underway. GSE is ready. We are getting telemetry.
Installation of instruments and cabling is underway. Function check of star camera is done, and one cable was found bad. Repaired. Instruments on base structure are mostly connected, and power-up tests will start tomorrow.
Our first crew arrived on Oct. 4. So far unpacking and assemble of structure have been done. Function check and installation of onboard instruments are underway. Pre-installation check of focal plane (position-sensitive NaI : HXI) has been done and found perfect. Mirrors are still in the container.
We got rent-a-car on Oct. 7. The hotel is in the middle of nowhere, and is about 4 km away from nearby town (Lorena).